OpenStudy (anonymous):

what is half of 2^{40} ?

OpenStudy (anonymous):

\[2^{40}\]

OpenStudy (anonymous):

\[2^{39}\]

OpenStudy (anonymous):

could you show me how you got that ?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

lets do a simpler one. \[2^5=32\] half of 32 is 16 \[2^4=16\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now for the "math" half of \[2^{40}\] means \[\frac{2^{40}}{2}\]

OpenStudy (anonymous):

when you divide you subtract the exponents, so you get \[2^{40-1}=2^{39}\]

OpenStudy (anonymous):

sorry im a little lost, why did u subtract 1 ?

OpenStudy (anonymous):

oh because the denominator was 2, and \[2=2^1\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

and when you divide you subtract the exponents. \[\frac{b^n}{b^m}=b^{n-m}\]

OpenStudy (anonymous):

but forget about that for a second. power of two look like 2,4,8,16,32,64,128,...

OpenStudy (anonymous):

multiply by 2 to get the next one. so if you divide by 2 you get the previous one.

OpenStudy (anonymous):

alright then thank you :)!

OpenStudy (anonymous):

o wow